the measure of an interior angle of a regular n gon is 140 degrees, how many sides are in the regular polygon?
thanks

Answer by stanbon(75887)

the measure of an interior angle of a regular n gon is 140 degrees, how many sides are in the regular polygon?
--------
The corresponding exterior angle must be 180-140 = 40 degrees
----
The sum of ALL the exterior angles is ALWAYS 360.
---
# of exterior angles = 360/40 = 9
Therefore, # of sides = 9
==================
Cheers,
Stan H.
===============